Calculate the energy of radiation emitted when an electron in a hydrogen atom makes each of the following transitions:

a) n = 3 to n = 1

b) n = 5 to n = 2
Answer #1

The energy of radiation emitted when an electron in a hydrogen atom makes transitions between is given by

E = -1312((1/n1^2)-(1/n2^2)) KJ

a) n = 3 to n = 1

The energy emitted, E = -1312((1/3^2)-(1/1)) = 1166.22 KJ

b). n = 5 to n = 2

The energy emitted, E = -1312((1/5^2)-(1/2^2)) = 275.52 KJ